OpenStudy (anonymous):

OKAY!! next one! Expand : \[\{ 3\frac { a }{ 2 } +\frac { b }{ 4 } -2c\} ^{ 2 }\\ \]

OpenStudy (anonymous):

\[\{ 3\frac { a }{ 2 } +\frac { b }{ 4 } -2c\} ^{ 2 }\\ \]

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

Expand : \[\{ 3\frac { a }{ 2 } +\frac { b }{ 4 } -2c\} ^{ 2 }\\ \]

mathslover (mathslover):

hint